Application of Newtons Laws: Two skaters, a man an
Answer #1

10 ans

step:1

mass of man m1=82 kg

mass of woman m2=48 kg

force F=45 N

step:2

now we find the acceleration of man

F=m1a1

45=82*a1

acceleration of man =a1=F/m1=45/82=0.55 m/s

the direction is east

step:3

now we find the magnitude of accelertion of woman

acceleration of woman a2=F/m2=45/48=0.9375 m/s^2

the direction is east